What is the first number if the second is 17 and the third is 3
Answer by Cromlix(4381)

Hi there,
a = first number
a + b = Second number
a + 2b = Third number
.......
a + b = 17.....(1)
a + 2b = 3.....(2)
Take (1) from (2)
....b = -14
Substitute b + -14
into (1)
a + b = 17
a + (-14) = 17
a - 14 = 17
a = 17 + 14
a = 31.
First number = 31
Hope this help:-)